OpenStudy (anonymous):

24. For a particular group of scores, M = 15 and SD = 3. Provide the raw score for a Z score of 7: (Points : 2) 15 24 36 48

OpenStudy (anonymous):

Use the formula, \[Z=\frac{X-\mu}{\sigma}\] Plug in the given values, \(Z=7\), \(\mu=M=15\), and \(\sigma=SD=3\).

OpenStudy (anonymous):

36??

OpenStudy (anonymous):

\[7=\frac{X-15}{3}~\Rightarrow~21=X-15~\Rightarrow~X=\cdots\]

OpenStudy (anonymous):

So yes, \(X=36\).